a women bought 7 ears of corn and 6 oranges for $2.44. a second women bought 10 ears of corn and 3 oranges for $2.26. find the price of an ear of corn and the price of an orange.
Answer by addingup(3677)

7c+6o= 2.44
10c+3o= 2.26
In the second equation multiply both sides times 2:
20c+6o= 4.52 Now subtract this equation from the first:
-------------------------------------
7c+6o= 2.44
-
20c+6o= 4.52
-------------
-13c= -2.08 Now divide both sides by -13:
c= 0.16 is the cost of an ear of corn. Now in the first equation let's use this value for c:
7(0.16)+6o= 2.44
1.12+6o= 2.44 subtract 1.12 on both sides:
6o= 1.32 Divide both sides by 6:
o= 0.22
-----------------------------------
Check. We just used the first equation to find o, now let's check to see if our answer makes the second equality true:
10c+3o= 2.26
10(0.16)+3(0.22)= 2.26
1.6+0.66= 2.26
2.26 = 2.26 We've got the correct answer